### v1

# Reducing Stochastic Games to Semidefinite Programming

Abstract
We present a polynomial-time reduction from max-average constraints to the feasibility problem for semidefinite programs. This shows that Condon’s simple stochastic games, stochastic mean payoff games, and in particular mean payoff games and parity games can all be reduced to semidefinite programming.
1 Introduction
There are two important clusters of computational problems that are not known to be in the complexity class P: the first is related to numeric computation, and contains the sums-of-square-roots problem [GGJ76, EHS24], the Euclidean shortest path problem, PosSLP [ABKPM09], and the feasibility problem for semidefinite programs [NN94, Ram97]. The second cluster is related to tropical geometry, and contains for instance the model checking problem for the propositional $\mu$-calculus [EJ91, EJS93], parity games [Mar75, Mos91], mean payoff games [EM79], and/or scheduling [MSS04], stochastic mean payoff games [AM09], and simple stochastic games [Con92]. So far, no polynomial-time reduction from a problem of one of the clusters to a problem from the other cluster was known. We show that all of the mentioned problems from the second cluster can be reduced in polynomial time to the feasibility problem for semidefinite programs from the first cluster.
Semidefinite programming is a generalisation of linear programming with many important algorithmic applications in discrete and continuous optimisation, both from a theoretical and a practical perspective [GLS94, BTN01, Las09]. However, already the feasibility problem for semidefinite programs is not known to be in P. By Ramana’s duality [Ram97], the problem is either in the intersection of NP and coNP, or outside of the union of NP and coNP. The semidefinite feasibility problem falls into the existential theory of the reals, which is known to be in PSPACE [Can88]. However, it has been observed by Khachiyan that the smallest feasible solution to an SDP might be of doubly exponential size (see, e.g., [PK97]), which is an obstacle for the polynomial-time algorithmic methods known for linear programming. In fact, it is possible to reduce the PosSLP problem to semidefinite programming [TV08]. PosSLP has been introduced in [ABKPM09], motivated by the ‘generic task of numerical computation’. For example, the sums-of-square-roots problem, which is a numeric computational problem not even known to be in NP and the barrier for polynomial-time tractability for numerous problems in computational geometry, such as the Euclidean shortest path problem, has a polynomial-time reduction to PosSLP [ABKPM09].
Mean payoff games are turn-based two-player games on graphs [EM79, ZP96]; it is a famous open problem in theoretical computer science whether there is a polynomial-time algorithm to decide for a given graph which of the players has a winning strategy. Finding the winning region in mean payoff games is polynomial-time equivalent to various other computational problems in theoretical computer science, for instance scheduling with and-or-constraints [MSS04], the max-atoms problem [BNRC08], as well as solvability of max-plus systems and tropical linear feasibility [ABG07, AGG12]. The latter can be seen as the tropical analog of testing feasibility of linear inequalities (which for classical geometry is known to be in P, e.g. via the ellipsoid method [Kha79, GLS94]).
Furthermore, there is a simple reduction from parity games to mean payoff games by using the priorities of the parity game with a suitably large basis [Pur95]. Parity games are polynomial-time equivalent to the model-checking problem of the propositional $\mu$-calculus [EJS93], which has been called ‘the most important logics in model checking’ [BW18]. For parity games, a quasipolynomial algorithm has been found recently [CJK+22]. Subsequent work indicates that the various quasipolynomial methods that have been obtained lately for parity games [LPSW22, JMT22, LB20] do not extend to mean payoff games [CFGO22].
[…]
Figure 1: Computational problems that are not known to be in P and not known to be NP-hard, and their relationship. Arcs indicate polynomial-time (many-one) reductions, the dotted arc indicates a polynomial-time Turing reduction.
We introduce a fragment of max-plus-average constraints, which we call max-average constraints. There is a polynomial-time reduction from stopping simple stochastic games to max-average constraints (Section 3). Max-average constraints have the advantage that they can be further reduced to the feasibility problem for semidefinite programs. This implies that all the mentioned computational problems from the second cluster can be reduced to semidefinite programming. See Figure 1 for an overview of the mentioned computational problems and their relationship.
The reduction uses a similar idea as the reduction of Schewe [Sch09] to reduce parity games and mean payoff games to linear programming, which is, however, not a proper polynomial-time reduction. The idea is to replace constraints of the form $x_{0}=\max(x_{1},\dots,x_{k})$ by $x_{0}=\log_{b}(b^{x_{1}}+\dots+b^{x_{k}})$ for a sufficiently large $b$. Following [AGS18], we extend this idea to max-average constraints and obtain a proper polynomial-time reduction to non-Archimedean SDPs. From there, we use bounds from quantifier elimination results to translate non-Archimedean SDPs to real SDPs. However, we still have the obstacle that these SDPs involve coefficients of doubly exponential size, which means that they have exponential representation size and do not lead to a polynomial-time reduction. We overcome this difficulty by using the duality of semidefinite programming to find small expressions that define coefficients of doubly exponential size, which combined with the above ideas leads to a polynomial-time reduction from max-average constraints to (real) SDPs.
Our results imply that if there is a polynomial-time algorithm for SDPs, then all the mentioned problems in the second cluster can be solved in polynomial time as well. Conversely, our reduction can be used to translate interesting families of instances for (stochastic) mean payoff games or parity games into SDP instances, which might yield interesting families of instances for (algorithmic approaches to) semidefinite programming. For linear programming, this idea turned out to be very fruitful: certain instances of parity games have been used to prove exponential lower bounds for several famous pivoting rules for the simplex method in linear programming [FHZ11, Fri11, FHZ14, AF17, DFH23]. However, these results were not based on a general polynomial-time reduction from parity games to linear programming (which is not known!). We also hope that algorithmic ideas for solving games might generalise to (certain classes of) SDPs.
Our results also show that two large research communities (convex real algebraic geometry, and more generally convex optimisation on the one hand, and verification and automata theory on the other hand) were working on closely related computational problems and might profit from joining forces in the future.
1.1 Related work
There are multiple works that reduce games to continuous optimization problems. As mentioned above, one-player mean payoff games (Markov decision processes) can be reduced to linear programming [Put05], which implies that they can be solved in polynomial time. For the case of two-player deterministic mean payoff games, Schewe [Sch09] proposed a reduction to linear programs with coefficients of exponential representation size. Such programs can also be interpreted as linear programs over non-Archimedean ordered fields [ABGJ15, ABGJ14, ABGJ21]. The idea of Schewe was generalized to stochastic games in [BEGM17], where the authors prove that stochastic mean payoff games can be encoded by convex programs with constraints of exponential encoding size. Such programs can be solved by the ellipsoid method in pseudo-polynomial time provided that the number of stochastic nodes is fixed. The non-Archimedean approach was generalized from deterministic games to stochastic games in the work [AGS18], which proves that stochastic mean payoff games can be reduced to semidefinite programs over non-Archimedean fields. The paper [AGS18] serves as a basis for our results.
These reductions express mean payoff games as convex optimization problems, but these problems are “non-standard” in the sense that they either use constraints of exponential encoding size or are expressed over non-Archimedean fields. There is yet another line of work deriving polynomial-time reductions from mean payoff games to classical optimization problems which are not convex. In particular, [Con93] expressed simple stochastic games as non-convex quadratic programs. This reduction is further studied in [KRSW22]. Another reduction to (generalized) complementarity problems with $P$-matrices was proposed in [GR05, SV06] and extended in subsequent works [JS08, FJS10, HIJ13]. Such linear complementarity problems form a subclass of quadratic programs and belong to the complexity class UniqueEOPL [FGMS20]. However, they are in general non-convex and can therefore not be expressed by semidefinite programs.
2 Preliminaries
The set of natural numbers $\{0,1,2,\dots\}$ is denoted by ${\mathbb{N}}$, and the set of real numbers by ${\mathbb{R}}$. We assume familiarity with first-order logic; see, e.g., [Hod97]. Let $\tau$ be a signature, i.e., a set of function and relation symbols. A first-order $\tau$-formula is a formula build from the Boolean connectives, atomic $\tau$-formulas and quantification in the usual way. If $A$ is a $\tau$-structure and $\phi$ is a $\tau$-formula with free variables $x_{1},\dots,x_{k}$, then $\phi$ defines over $A$ the relation
$\{(a_{1},\dots,a_{k})\in A^{k}\mid A\models\phi(a_{1},\dots,a_{k})\}.$
A set if called semialgebraic if it has a first-order definition over the field $({\mathbb{R}};+,\cdot)$, allowing real parameters. A first-order formula $\phi(x_{1},\dots,x_{k})$ with free variables $x_{1},\dots,x_{k}$ is called primitive positive if it is of the form
$\exists y_{1},\dots,y_{\ell}(\psi_{1}\wedge\cdots\wedge\psi_{m})$
where $\psi_{1},\dots,\psi_{m}$ are atomic formulas over the variables $x_{1},\dots,x_{k},y_{1},\dots,y_{\ell}$. A relation is called primitive positive definable in a $\tau$-structure $A$ if there exists a primitive positive $\tau$-formula which defines the relation over $A$.
2.1 Semidefinite Programming
For a symmetric matrix $S\in{\mathbb{R}}^{m\times m}$ we write $S\succeq 0$ if $S$ is positive semidefinite, i.e., if $x^{\top}Sx\geq 0$ for all $x\in{\mathbb{R}}^{m}$, and $S\succ 0$ if it is positive definite, i.e., if $x^{\top}Sx>0$ for all $x\in{\mathbb{R}}^{m}\setminus\{0\}$. A semidefinite program (SDP) is an optimisation problem of the following form:
$\inf$ $\sum_{i=1}^{n}c_{i}x_{i}$
such that $\sum_{i=1}^{n}x_{i}A_{i}-B\succeq 0\,,$
where $A_{1},\dots,A_{n},B\in{\mathbb{R}}^{m\times m}$ are symmetric and $c_{1},\dots,c_{n}\in{\mathbb{R}}$. If the data $A_{1},\dots,A_{n},B,c_{1},\dots,c_{n}$ is rational, then the problem of deciding whether the supremum exists, and if so, to compute the supremum, is a computational problem known as the semidefinite programming problem. The representation size of the SDP then consists of the sum of the bit lengths of all the rational numbers in the input.
The expression $\sum_{i=1}^{n}x_{i}A_{i}-B\succeq 0$ in Section 2.1 is called a linear matrix inequality (LMI), and defines the feasible region of the SDP, which is called a spectrahedron. The representation size of LMIs is defined analogously as for SDPs. We say that an SDP is strictly feasible if $\sum_{i=1}^{n}x_{i}A_{i}-B\succ 0$ has a solution $x\in{\mathbb{R}}^{n}$.
Definition 2.1.
Given symmetric matrices $A_{1},\dots,A_{n},B\in{\mathbb{Q}}^{m\times m}$ the feasibility problem for semidefinite programs asks to decide whether there is a solution $x\in\mathbb{R}^{n}$ to an LMI
$x_{1}A_{1}+\dots+x_{n}A_{n}-B\succeq 0\enspace.$
Note that spectrahedra are convex and semialgebraic. It has been conjectured that every convex semialgebraic set is the projection of a spectrahedron (also called a spectrahedral shadow), but this turned out to be false [Sch18] (see [BKT24] for an alternative proof of an even stronger inexpressibility result). It is easy to see that the set of spectrahedral shadows is closed under primitive positive definability, and that for every primitive positive formula $\phi$ we may find an LMI for the spectrahedral shadow defined by $\phi$ whose representation size is bounded by the sum of the representation sizes of conjuncts of $\phi$. We also note that the definition of a semidefinite program makes sense over any real closed field, not only over the real numbers. In particular, it is meaningful to consider semidefinite programs over non-Archimedean real closed fields, such as the field of Puiseux series which we discuss in Section 4.
The (Frobenius) inner product of two matrices $S,T\in{\mathbb{R}}^{m\times m}$ is defined as $\langle S,T\rangle:=\text{trace}(ST)=\sum_{i,j}S_{ij}T_{ij}$. The dual program to the SDP from Section 2.1 (to which we refer as the primal) is given by
$\sup_{Y\succeq 0}$ $\langle B,Y\rangle$ (D)
such that $\langle A_{i},Y\rangle=c_{i}\text{ for every }i\in\{1,\dots,n\}\,.$
It is well-known that the value of the primal SDP is bounded from below by the value of the dual; however, the lower bound might not be tight, i.e., there might be a non-zero duality gap. Moreover, both primal and dual might not have optimal solutions, i.e., the supremum may not be a maximum and the infimum might not be a minimum. See [Ram97] for a more complicated, but exact dual. However, for us the following sufficient condition for a vanishing duality gap will be enough.
Theorem 2.2 (see Theorem 2.4.1 in [BTN01]).
If the primal SDP is strictly feasible and bounded from below, then the dual program attains its supremum and there is no duality gap.
2.2 Stochastic Games
A stochastic mean payoff game is a two-player game played on a finite edge-weighted directed graph $\vec{\mathcal{G}}=(V,E,r)$, i.e., $V$ is a finite set of nodes, $E\subseteq V^{2}$ is a set of edges, and $r\colon E\to{\mathbb{Q}}$ is a weight function. The nodes of the graph are partitioned into three subsets, $V=V_{\mathrm{Min}}\sqcup V_{\mathrm{Max}}\sqcup V_{\mathrm{Rand}}$. The nodes in $V_{\mathrm{Min}}$ are controlled by player Min, the nodes in $V_{\mathrm{Max}}$ are controlled by player Max, and the nodes in $V_{\mathrm{Rand}}$ are controlled by nature. The players play the game by moving a pawn on the graph. When the pawn is on a node $v\in V_{\mathrm{Min}}$, player Min chooses an outgoing edge $(v,u)$ and moves the pawn to $u$. Analogously, player Max chooses the next move when the pawn is on a node $v\in V_{\mathrm{Max}}$. When the pawn is on a node $v\in V_{\mathrm{Rand}}$, the next move is chosen by nature randomly, according to some probability distribution on the outgoing edges of $v$. This probability distribution is known to both players and does not change during the game. To ensure that players can always make a move, we suppose that all nodes of the graph have at least one outgoing edge. The weights of the edges of the graph represent payoffs that player Max receives from player Min after each move of the pawn. In order to state the objectives of both players, we first specify the strategies that they are allowed to use. A pure and positional strategy (also called a policy) for player Min is a function $\sigma\colon V_{\mathrm{Min}}\to V$ that to each node controlled by Min associates a neighbor, $(v,\sigma(v))\in E$. Player Min plays according to $\sigma$ if they always move the pawn to $\sigma(v)$ when the pawn lands on $v$. We analogously define a policy $\tau\colon V_{\mathrm{Max}}\to V$ of player Max. Note that if we fix a pair of policies and an initial position of the pawn, then the resulting movement of the pawn is a Markov chain on the graph $(V,E)$. Given an initial position $u\in V$ of the pawn and a pair of policies $(\sigma,\tau)$, the mean payoff reward of player Max is the quantity
$g(\sigma,\tau)_{u}\coloneqq\lim_{N\to\infty}\frac{1}{N}\mathbb{E}^{u}_{\sigma,\tau}(r_{u_{0}u_{1}}+\dots+r_{u_{N-1}u_{N}})\,,$
where $u=u_{0},u_{1},\dots,u_{N}$ is the path of the pawn in the first $N$ turns of the game and the expectation is taken with respect to the Markov chain obtained by fixing $(\sigma,\tau)$. The ergodic theorem of Markov chains shows that $g(\sigma,\tau)_{u}$ is well-defined. The objective of player Max is to maximize the mean payoff reward while the objective of player Min is to minimize this reward. It is known that both players in a stochastic mean payoff game have optimal policies [LL69]. In other words, there exist policies $(\sigma^{*},\tau^{*})$ and a vector $\chi\in\mathbb{R}^{V}$ such that for every $u$ and every $(\sigma,\tau)$ we have the inequality
$g(\sigma^{*},\tau)_{u}\leq\chi_{u}\leq g(\sigma,\tau^{*})_{u}\,.$
The reward $\chi=g(\sigma^{*},\tau^{*})$ is called the value of the game. In order to make these games computationally tractable, we suppose that the weights and the probability distributions at nodes in $V_{\mathrm{Rand}}$ are rational. The nodes of the graph are also called the states of the game.
A simple stochastic game is a special case of a stochastic mean payoff game. In this game, every node in $V_{\mathrm{Rand}}$ has out-degree exactly two and the probability distribution at every such node is $(1/2,1/2)$. Furthermore, the graph is equipped with two special terminal states called $\mathrm{Win}$ and $\mathrm{Lose}$. These states are absorbing, i.e., $\mathrm{Win}$ has only one outgoing edge that goes back to $\mathrm{Win}$ and $\mathrm{Lose}$ has only one outgoing edge that goes back to $\mathrm{Lose}$. The weight of the edge going from $\mathrm{Win}$ to $\mathrm{Win}$ is equal to $1$, while the weight of all of the other edges of the graph is equal to 00. In this case, the mean payoff reward of player Max is simply the probability that the pawn reaches the winning state, $g(\sigma,\tau)_{u}=\mathbb{P}^{u}_{\sigma,\tau}(\text{Reach}(\mathrm{Win}))$. A simple stochastic game is called stopping if for every initial position $u\in V$ and every pair of policies $(\sigma,\tau)$, the pawn reaches a terminal state with probability one, $\mathbb{P}^{u}_{\sigma,\tau}(\text{Reach}(\mathrm{Win},\mathrm{Lose}))=1$. In other words, the game is stopping if the Markov chain obtained by fixing any pair of policies has only two recurrent classes, namely $\{\mathrm{Win}\}$ and $\{\mathrm{Lose}\}$.
The computational problem associated with stochastic mean payoff games is the following: given a stochastic mean payoff game, find its value and a pair of optimal policies. In [AM09], Andersson and Miltersen proved that this problem is polynomial-time (Turing) reducible to the problem of finding the value of a stopping simple stochastic game. By a binary search argument, the problem of computing the value of a stopping simple stochastic game can be further reduced to the decision problem, in which we want to decide whether the value of a given state of such a game is at least $1/2$. The details of this reduction can be found, e.g., in [AGS18] or [Sko18, Section 7.1].
Theorem 2.3 ([AGS18, Lemma 48]).
The problem of computing the value and a pair of optimal policies of a stochastic mean payoff game is polynomial-time (Turing) reducible to the problem of deciding whether a given state of a stopping simple stochastic game has value at least $1/2$.
3 From stochastic games to max-average constraints
In this section, we focus on the decision problem stated in Theorem 2.3 above: given a stopping simple stochastic game we want to decide whether the value of a given state is at least $1/2$. This problem can be reduced to max-average constraint satisfaction problems. The ideas for this reduction are not new, but we are not aware of an explicit reference for this fact, and we therefore present a short proof in this section. In the next section, we will show that max-average CSPs then further reduce to the feasibility problem of non-Archimedean semidefinite programs. The fact that stochastic games can be encoded as non-Archimedean SDPs was proven in [AGS18] and our reduction is very similar. While their approach relies on reducing stochastic games to ergodic games, we omit this step and obtain a slightly more straightforward reduction.
Definition 3.1.
An instance of a max-average constraint satisfaction problem is a conjunction of constraints, each of which has one of the following forms:
  1. 1.
    $x_{0}\leq\max(x_{1},\dots,x_{k})$ where $k\in\mathbb{N}$,
  2. 2.
    $x_{0}\leq\frac{x_{1}+x_{2}}{2}$,
  3. 3.
    $x_{0}=c$ where $c$ is a rational constant.
The computational task is to determine whether such a conjunction has a solution over ${}\mathbb{Q}\cup\{-\infty\}$ (equivalently, over ${}\mathbb{R}\cup\{-\infty\}$).
Note that max-average CSP instances may also use constraints of the form $x_{0}\leq\min(x_{1},\dots,x_{k})$ where $k\in\mathbb{N}$, $x_{0}\leq c$, or $x_{0}\geq c$ for some rational constant $c$, because the respective relations are primitively positively definable over the other relations.
We now show that the problem of solving stochastic mean payoff games can be reduced to max-average CSPs. A reduction to slightly more general CSPs, including constraints of the form $x_{0}\leq c+x_{1}$, follows from [AGS18, Theorem 18] and is based on a Collatz–Wielandt property of order-preserving and additively homogeneous maps. However, for our purposes, it is useful that the reduction does not use these additional constraints. Instead, we give a proposition that characterizes the value of a stopping simple stochastic game. It was already proven in [Con92] that this value is a unique fixed point of an operator involving min, max, and averaging operations. By replacing the fixed point equations with inequalities, we obtain a set of certificates which prove that the value is at least $1/2$. This observation is proven, e.g., in [BBDFS23] and gives inequalities very similar to the inequalities on Shapley operators used in [AGS18].
Proposition 3.2.
Consider a stopping simple stochastic game played on a directed graph $\vec{\mathcal{G}}=([n],E)$. Then, the value of a state $k\in[n]$ is at least $1/2$ if and only if the following system of inequalities (an instance of a max-average constraint satisfaction problem) has a solution $x\in\mathbb{Q}_{\geq 0}^{n}$:
$x_{i}$ $\leq\max_{(i,j)\in E}\{x_{j}\}$ for every $\text{for every }\in V_{\mathrm{Max}},$ (1)
$x_{i}$ $\leq\min_{(i,j)\in E}\{x_{j}\}$ for every $\text{for every }i\in V_{\mathrm{Min}},$
$x_{i}$ $\leq\frac{1}{2}\sum_{(i,j)\in E}x_{j}$ for every $\text{for every }i\in V_{\mathrm{Rand}},$
$x_{\mathrm{Lose}}$ $=0,$
$x_{\mathrm{Win}}$ $=1,$
$x_{k}$ $\geq 1/2.$
Proof.
Let $(Q)$ denote the system of inequalities obtained from Eq. 1 by removing the inequality $x_{k}\geq 1/2$. By [Con92] or [BBDFS23, Theorem 77], the value of the game satisfies $(Q)$ with inequalities replaced by equalities. Hence, if the value at state $k$ is at least $1/2$, then the value of the game satisfies Eq. 1. Moreover, the value of the game is a rational vector by [Con92, Lemma 2] or [BBDFS23, Corollary 18]. Conversely, if a point $x\in\mathbb{Q}_{\geq 0}^{n}$ satisfies $(Q)$, then [BBDFS23, Lemma 83] implies that for all $i\in[n]$, the entry $x_{i}$ is not greater than the value of the game at state $i$. Hence, if we further suppose that $x_{k}\geq 1/2$, then the value of the game at state $k$ is at least $1/2$. ∎
4 From max-average constraints to non-Archimedean SDPs
By “lifting” the inequalities Eq. 1 to Puiseux series, we get the reduction from max-average constraint satisfaction problems to non-Archimedean SDPs. A (formal real) Puiseux series is a series of the form
$\bm{x}=\bm{x}(t)=c_{0}t^{a/n}+c_{1}t^{(a-1)/n}+c_{2}t^{(a-2)/n}+\dots\,,$
where $t$ is a formal parameter, the number $n$ is a positive integer, $a$ is an integer, and the coefficients $c_{i}$ are real. We further assume that $c_{0}\neq 0$ for every Puiseux series except the zero series. The number $a/n$ is called the valuation of the series $\bm{x}$ and is denoted by $\operatorname*{\mathsf{val}}(\bm{x})$, with the convention that $\operatorname*{\mathsf{val}}(0)=-\infty$. Puiseux series can be added and multiplied in the natural way. Moreover, they can be ordered by defining $\bm{x}\geq 0$ if $c_{0}\geq 0$ and $\bm{x}\geq\bm{y}$ if $\bm{x}-\bm{y}\geq 0$. It is known that Puiseux series form a real closed field, see, e.g., [BPR06, Section 2.6] for a detailed discussion. [^1] In this work, we use a convention in which one interprets $t$ as a “very large” parameter. Many authors use the opposite but equivalent convention in which $t$ is a very small parameter. This explains the differences in notation between our work and [BPR06]. We denote the field of Puiseux series by $\mathbb{K}$. For any $\bm{x},\bm{y}\in\mathbb{K}$, the following properties of the valuation function are immediate from the definitions:
$\operatorname*{\mathsf{val}}(\bm{x}+\bm{y})$ $\leq\max\{\operatorname*{\mathsf{val}}(\bm{x}),\operatorname*{\mathsf{val}}(\bm{y})\}\,,$ (2)
$\operatorname*{\mathsf{val}}(\bm{x}\bm{y})$ $=\operatorname*{\mathsf{val}}(\bm{x})+\operatorname*{\mathsf{val}}(\bm{y})\,,$
$\bm{x}\geq\bm{y}\geq 0$ $\implies\operatorname*{\mathsf{val}}(\bm{x})\geq\operatorname*{\mathsf{val}}(\bm{y})\,.$
Theorem 4.1.
An instance $\phi$ with variables $x_{1},\dots,x_{n}$ of a max-average constraint satisfaction problem has a solution in ({}$(\mathbb{Q}\cup\{-\infty\})^{n}$ if and only if the following system of inequalities $S_{\phi}(t)$, which is a (quadratic) semidefinite program over Puiseux series, has a solution in $\mathbb{K}^{n}$: for every variable $x_{i}$ of $\phi$ we have a variable $\bm{x}_{i}$, and the inequalities are
$\bm{x}_{i_{0}}$ $\leq\bm{x}_{i_{1}}+\cdots+\bm{x}_{i_{k}}$ for every constraint $x_{i_{0}}\leq\max(x_{i_{1}},\dots,x_{i_{k}})$ in $\phi$, (3)
$\bm{x}^{2}_{i_{0}}$ $\leq\bm{x}_{i_{1}}\bm{x}_{i_{2}}$ for every constraint $x_{i_{0}}\leq\frac{x_{i_{1}}+x_{i_{2}}}{2}$ in $\phi$,
$\bm{x}_{i_{0}}$ $=t^{c}$ for every constraint $x_{i_{0}}=c$ in $\phi$,
$\bm{x}_{i_{0}}$ $\geq 0$ for every $i_{0}\in\{1,\dots,n\}$.
Proof.
Suppose that $\bm{x}=(\bm{x}_{1},\dots,\bm{x}_{n})\in\mathbb{K}^{n}$ satisfies Eq. 3. Then, $\operatorname*{\mathsf{val}}(\bm{x})$ belongs to ({}$(\mathbb{Q}\cup\{-\infty\})^{n}$ and satisfies $\phi$ by the properties of the valuation function Eq. 2. Conversely, if $(x_{1},\dots,x_{n})\in(\mathbb{Q}\cup\{-\infty\})^{n}$ satisfies $\phi$, then $(t^{x_{1}},\dots,t^{x_{n}})$ (with the convention that $t^{-\infty}=0$) satisfies Eq. 3.
This amounts to the feasibility problem for a semidefinite program: the linear (in)equalities can be expressed by diagonal matrices and each quadratic condition in combination with the non-negativity of all variables is just the positive semidefiniteness of a symmetric 2-by-2 matrix. Conjunctions of such constraints can be expressed by LMIs for block diagonal matrices. ∎
5 From non-Archimedean to real SDPs
To conclude, we aim to reduce the non-Archimedean semidefinite program from Eq. 3 to a real semidefinite program. We will do this in two steps. First, we replace the formal parameter $t$ by a real number that is sufficiently large (doubly exponential in the size of the CSP). In fact, any non-Archimedean semialgebraic feasiblity problem can be reduced to a semialgebraic feasibility problem over the reals by replacing $t$ with a sufficiently large real number. This follows from the quantifier elimination in real closed fields coupled with the definition of the order in Puiseux series. By using the effective bounds on quantifier elimination discussed, e.g., in [BPR06], we can give a bound on this “sufficiently large” number. Second, we will show that this SDP with very large coefficients can be reduced to another SDP that has coefficients of polynomial bit-size.
Before presenting the proofs, we note that a reduction from non-Archimedean SDPs coming from games to real SDPs with large coefficients was already studied in [AGS18, Theorem 36]. However, the bounds given in [AGS18] rely on an additional assumption on the underlying game. In particular, they cannot be applied directly to the SDP given in Eq. 3. For max-average systems coming from games, one could overcome this restriction by perturbing the underlying game slightly. Nevertheless, the bounds coming from quantifier elimination which we present below do not require any additional assumptions and can be of independent interest.
Theorem 5.1.
Let $P_{1},\dots,P_{m}\in\mathbb{Z}[x_{0},x_{1},\dots,x_{n}]$ be a collection of polynomials with integer coefficients and let {,,<,>,=,}$\diamond\in\{\leq,\geq,<,>,=,\neq\}^{m}$ be a sign pattern. Consider the semialgebraic set over Puiseux series defined by
$\bm{S}\coloneqq\{\bm{x}\in\mathbb{K}^{n}\colon\forall i\in[m],\,P_{i}(t,\bm{x})\diamond_{i}0\}\,.$
Furthermore, for any real $L>0$ consider the semialgebraic set over the real numbers defined by
$S(L)\coloneqq\{x\in\mathbb{R}^{n}\colon\forall i\in[m],\,P_{i}(L,x)\diamond_{i}0\}\,.$
Then, $\bm{S}$ is nonempty if and only if $S(L)$ is nonempty for all sufficiently large $L$. More precisely, there exists an absolute constant $C>0$ with the following property. If the maximal degree of any polynomial $P_{i}$ is bounded by $d\geq 2$ and $s$ denotes the maximal bit-size of any coefficient of $P_{i}$, then for $L>2^{sd^{Cn}}$ either all $S(L)$ are empty or all $S(L)$ are nonempty. In the former case $\bm{S}$ is empty and in the latter case $\bm{S}$ is nonempty.
Proof.
Consider the first-order formula $\phi(x_{0})$ in the language of ordered rings given by
$\exists x_{1},\dots,\exists x_{n}\operatorname*{\bigwedge}_{i=1}^{m}(P_{i}(x_{0},x)\diamond_{i}0)\,.$
By Tarski’s quantifier elimination theorem [BPR06, Theorem 2.77], there exists a quantifier-free formula $\psi(x_{0})$ that is equivalent to $\phi(x_{0})$ in the theory of real closed fields. The bounds on quantifier elimination given in [BPR06, Theorem 14.6] show that $\psi(x_{0})$ can be taken to be of the form
$\operatorname*{\bigvee}_{i=1}^{I}\operatorname*{\bigwedge}_{j=1}^{J_{i}}\left(\operatorname*{\bigvee}_{k=1}^{N_{ij}}\operatorname*{\mathrm{sign}}(Q_{ijk}(x_{0}))=\delta_{ijk}\right)\,,$
where $\delta_{ijk}\in\{-1,0,1\}$ and $Q_{ijk}\in\mathbb{Z}[x_{0}]$ are univariate polynomials with integer coefficients. Furthermore, every $Q_{ijk}$ has degree bounded by $d^{O(n)}$ and every coefficient of every $Q_{ijk}$ has bit-size bounded by $sd^{O(n)}$. Hence, the Cauchy bound [BPR06, Corollary 10.4] implies that every real root of $Q_{ijk}$ has absolute value bounded by $2^{sd^{O(n)}}$ (provided that $Q_{ijk}$ is a non-zero polynomial). In particular, the sign of $Q_{ijk}(L)$ is the same for all $L>2^{sd^{O(n)}}$. Therefore, either $S(L)$ are empty for all such $L$ or they are nonempty for all such $L$. Moreover, $Q_{ijk}(t)\in\mathbb{K}$ is a Puiseux series with finitely many terms. By definition, the sign of this series is the same as the sign of $Q_{ijk}(L)\in\mathbb{R}$. Hence, $\bm{S}$ is nonempty if and only if $S(L)$ is nonempty. ∎
In the next proposition and later, we use the following notation. Fix an instance $\phi$ of a max-average CSP and let $(c_{i})_{i}$ be the rational constants used in the constraints of $\phi$. For every $i$ we denote $c_{i}=p_{i}/q_{i}$ where $p_{i}\in\mathbb{Z}$ and $q_{i}\in\mathbb{N}^{*}$ and we put $W\coloneqq 2+\max_{i}\{|p_{i}|+q_{i}\}$.
Proposition 5.2.
There exists an absolute integer constant $M>0$ such that an instance $\phi$ of a max-average constraint satisfaction problem has a solution if and only if the real SDP $S_{\phi}(L)$ obtained from $S_{\phi}(t)$ (see Theorem 4.1) by replacing $t$ with $L>2^{2^{Mn\log_{2}W}}$, is feasible.
Proof.
Consider the system of inequalities given in Eq. 3. Since $\bm{x}_{i}\geq 0$ for all $i$, any equality of the form $\bm{x}_{i}=t^{c_{j}}=t^{p_{j}/q_{j}}$ can be replaced with an equivalent equality $\bm{x}_{i}^{q_{j}}=t^{p_{j}}$ if $p_{j}\geq 0$ and $t^{-p_{j}}\bm{x}_{i}^{q_{j}}=1$ if $p_{j}<0$. If we further replace the parameter $t$ by a variable symbol $x_{0}$, Eq. 3 becomes a system of polynomial inequalities of degree at most $W$ and such that the coefficients of the polynomials appearing in these inequalities are integers in $\{-1,0,1\}$. Hence, by Theorem 5.1, there exists $M>0$ such that the system Eq. 3 has a solution over Puiseux series if and only if the system obtained by replacing $t$ with $L>2^{2^{Mn\log_{2}W}}$ has a solution over the real numbers. Therefore, the claim follows from Theorem 4.1. ∎
In Proposition 5.2, we obtain a (real) semidefinite program, whose size, however, is exponential in the size of $\phi$ and therefore does not provide a polynomial-time reduction to the semidefinite program feasiblity problem. However, by the following result we can also reduce to a semidefinite program of polynomial representation size.
Theorem 5.3.
For every $n\in{\mathbb{Z}}$, the singleton set $\{2^{n}\}$ is the projection of a spectrahedron whose representation size is polynomial in the bit-size of $n$.
Proof.
To prove the claim, we extend the constructions of SDPs with solutions of doubly-exponential size given in [PK97, Ram97]. For $n=0$, the set $\{x\in\mathbb{R}\mid x=1\}$ is the desired spectrahedron. Suppose that $n\neq 0$ and let $|n|=\sum_{i=0}^{k}\bar{b}_{k-i}2^{i}$ be the binary representation of $|n|$ for $\bar{b}_{i}\in\{0,1\}$. Let $b_{i}=\operatorname*{\mathrm{sign}}(n)\bar{b}_{i}\in\{-1,0,1\}$ for all $i$, so that $n=\sum_{i=0}^{k}b_{k-i}2^{i}$. We first show that the set $\{x\in{\mathbb{R}}\mid x\geq 2^{n}\}$ is the projection of a spectrahedron $S$ of representation size that is polynomial in the bit-size of $n$. Let $S$ be defined by the linear matrix inequality
$A_{0}x_{0}+\cdots+A_{k}x_{k}-B=\begin{pmatrix}x_{0}&1&\\ 1&2^{-b_{0}}\\ &&x_{1}&x_{0}\\ &&x_{0}&2^{-b_{1}}\\ &&&&\ddots\\ &&&&&x_{k}&x_{k-1}\\ &&&&&x_{k-1}&2^{-b_{k}}\end{pmatrix}\succeq 0\,.$
This LMI expresses that
$x_{0}\geq 2^{b_{0}},\,x_{1}\geq 2^{b_{1}}x_{0}^{2},\,\dots,\,x_{k}\geq 2^{b_{k}}x_{k-1}^{2}\,.$
By induction, for every $\ell\geq 0$ we get $x_{\ell}\geq 2^{b_{0}2^{\ell}+b_{1}2^{\ell-1}+\dots+b_{\ell}2^{0}}$. Moreover, the vector $(x_{0},\dots,x_{k})$ with $x_{\ell}=2^{b_{0}2^{\ell}+b_{1}2^{\ell-1}+\dots+b_{\ell}2^{0}}$ for every $\ell\in\{0,\dots,k\}$ satisfies this LMI, so we obtain the desired set by projecting $S$ to the last coordinate. We use strong duality to obtain the singleton set $\{2^{n}\}$ as follows. Consider the SDP
$\inf\,(0\cdot x_{0}+\cdots+0\cdot x_{k-1}+1\cdot x_{k})$
over the spectrahedron $S$ specified above. Then, the dual is of the form $\sup\,\langle B,Y\rangle$ such that $\langle A_{i},Y\rangle=c_{i}$ for every $i\in\{0,\dots,k\}$ and $Y\succeq 0$, where $(c_{0},\dots,c_{k})=(0,\dots,0,1)$. Clearly, the primal is strictly feasible and bounded by $2^{n}$ from below, so we may use strong duality (Theorem 2.2) and obtain that $\sup\,\langle B,Y\rangle=2^{n}$ and that the supremum is attained for some $Y$. Then the following primitive positive formula over spectrahedral sets with one free variable $x_{k}$ defines the singleton relation $\{2^{n}\}$:
$\exists x_{0},\dots,x_{k-1},y_{1,1},\dots,y_{m,m}:A_{0}x_{0}+\cdots+A_{k}-B\succeq 0$
$\wedge\,\langle A_{0},Y\rangle=0\,\wedge\cdots\wedge\,\langle A_{k-1},Y\rangle=0\,\wedge\,\langle A_{k},Y\rangle=1$ $\wedge Y\succeq 0$
$\wedge\langle B,Y\rangle=x_{k}$
where $Y$ denotes $(y_{i,j})_{i,j\in\{1,\dots,m\}}$ for $m=2(k+1)$. As we have mentioned before, spectrahedral shadows are closed under primitive positive definability, and the representation size of the resulting LMI is bounded by the representation sizes for the conjuncts. We thus obtain an LMI of representation size that is polynomial in $k$. ∎
Corollary 5.4.
There is a polynomial-time reduction from the max-average constraint satisfaction problem to the feasibility problem for semidefinite programs.
Proof.
We have already specified a first reduction to a (real) SDP in Proposition 5.2. To obtain a proper polynomial-time reduction, we modify this SDP as follows. Fix an instance $\phi$; let $c_{i}$, $p_{i}$, $q_{i}$, and $W$ be as introduced before Proposition 5.2. Define $D\coloneqq\prod_{i}q_{i}$ and $K\coloneqq 2DW^{Mn}=D2^{1+Mn\log_{2}W}$, where $M$ is the integer constant from Proposition 5.2. Furthermore, let $L\coloneqq 2^{K}$. By Proposition 5.2, $\phi$ has a solution if and only if $S_{\phi}(L)$ has a solution over the real numbers. This SDP has constraints of the form $x_{i}=L^{c_{j}}=2^{c_{j}K}$. Since $K$ is an integer divisible by $D$, the number $c_{j}K$ is an integer for all $j$. Hence, by introducing new variables, we may replace the constraints $x_{j}=2^{c_{j}K}$ by the corresponding LMIs of small representation size from Theorem 5.3. The resulting SDP can be computed in polynomial time. Furthermore, it is feasible if and only if the original SDP is feasible, which proves the statement. ∎
Corollary 5.5.
There is a polynomial-time reduction from stopping simple stochastic games to the feasibility problem for semidefinite programs.
Proof.
Combine Corollary 5.4 with Proposition 3.2. ∎
Acknowledgments
This research benefited from the support of the FMJH Program PGMO. Manuel Bodirsky received funding from the ERC (Grant Agreement no. 101071674, POCOCOP). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency.

### v2

@@
-Reducing Stochastic Games to Semidefinite Programming
+Reducing Stochastic Games to Semidefinite Program Feasibility
 Abstract
-We present a polynomial-time reduction from max-average constraints to the feasibility problem for semidefinite programs.
+We present a polynomial-time reduction from max-plus-average constraints to the feasibility problem for semidefinite programs.
 This shows that Condon’s simple stochastic games, stochastic mean payoff games, and in particular mean payoff games and parity games can all be reduced to semidefinite programming.
+keywords:
+Mean-payoff games, stochastic games, semidefinite programming, max-plus-average constraints, max-atom problem
 1 Introduction
-There are two important clusters of computational problems that are not known to be in the complexity class P: the first is related to numeric computation, and contains the sums-of-square-roots problem [GGJ76, EHS24], the Euclidean shortest path problem, PosSLP [ABKPM09], and the feasibility problem for semidefinite programs [NN94, Ram97].
-The second cluster is related to tropical geometry, and contains for instance the model checking problem for the propositional $\mu$-calculus [EJ91, EJS93], parity games [Mar75, Mos91], mean payoff games [EM79], and/or scheduling [MSS04], stochastic mean payoff games [AM09], and simple stochastic games [Con92].
+There are two important clusters of computational problems that are not known to be in the complexity class P: the first is related to numeric computation, and contains the sums-of-square-roots problem [35, 27], the Euclidean shortest path problem, PosSLP [7], and the feasibility problem for semidefinite programs [51, 55].
+The second cluster is related to tropical geometry, and contains for instance the model checking problem for the propositional $\mu$-calculus [29, 28], parity games [48, 50], mean payoff games [26], and/or scheduling [49], stochastic mean payoff games [8], simple stochastic games [23], tropically convex and max-plus-average constraints [17].
 So far, no polynomial-time reduction from a problem of one of the clusters to a problem from the other cluster was known.
 We show that all of the mentioned problems from the second cluster can be reduced in polynomial time to the feasibility problem for semidefinite programs from the first cluster.
-Semidefinite programming is a generalisation of linear programming with many important algorithmic applications in discrete and continuous optimisation, both from a theoretical and a practical perspective [GLS94, BTN01, Las09].
+Semidefinite programming is a generalisation of linear programming with many important algorithmic applications in discrete and continuous optimisation, both from a theoretical and a practical perspective [37, 12, 44].
 However, already the feasibility problem for semidefinite programs is not known to be in P.
-By Ramana’s duality [Ram97], the problem is either in the intersection of NP and coNP, or outside of the union of NP and coNP.
-The semidefinite feasibility problem falls into the existential theory of the reals, which is known to be in PSPACE [Can88].
-However, it has been observed by Khachiyan that the smallest feasible solution to an SDP might be of doubly exponential size (see, e.g., [PK97]), which is an obstacle for the polynomial-time algorithmic methods known for linear programming.
-In fact, it is possible to reduce the PosSLP problem to semidefinite programming [TV08].
-PosSLP has been introduced in [ABKPM09], motivated by the ‘generic task of numerical computation’.
-For example, the sums-of-square-roots problem, which is a numeric computational problem not even known to be in NP and the barrier for polynomial-time tractability for numerous problems in computational geometry, such as the Euclidean shortest path problem, has a polynomial-time reduction to PosSLP [ABKPM09].
-Mean payoff games are turn-based two-player games on graphs [EM79, ZP96]; it is a famous open problem in theoretical computer science whether there is a polynomial-time algorithm to decide for a given graph which of the players has a winning strategy.
-Finding the winning region in mean payoff games is polynomial-time equivalent to various other computational problems in theoretical computer science, for instance scheduling with and-or-constraints [MSS04], the max-atoms problem [BNRC08], as well as solvability of max-plus systems and tropical linear feasibility [ABG07, AGG12].
-The latter can be seen as the tropical analog of testing feasibility of linear inequalities (which for classical geometry is known to be in P, e.g.
-via the ellipsoid method [Kha79, GLS94]).
-Furthermore, there is a simple reduction from parity games to mean payoff games by using the priorities of the parity game with a suitably large basis [Pur95].
-Parity games are polynomial-time equivalent to the model-checking problem of the propositional $\mu$-calculus [EJS93], which has been called ‘the most important logics in model checking’ [BW18].
-For parity games, a quasipolynomial algorithm has been found recently [CJK+22].
-Subsequent work indicates that the various quasipolynomial methods that have been obtained lately for parity games [LPSW22, JMT22, LB20] do not extend to mean payoff games [CFGO22].
-[…]
+By Ramana’s duality [55], the problem is either in the intersection of NP and coNP, or outside of the union of NP and coNP.
+The semidefinite feasibility problem falls into the existential theory of the reals, which is known to be in PSPACE [21].
+However, it has been observed by Khachiyan that the smallest feasible solution to an SDP might be of doubly exponential size (see, e.g., [52]), which is an obstacle for the polynomial-time algorithmic methods known for linear programming.
+In fact, it is possible to reduce the PosSLP problem (testing positivity of a value computed by a Straight Line Program) to semidefinite programming [60].
+PosSLP has been introduced in [7], motivated by the ‘generic task of numerical computation’.
+For example, the sums-of-square-roots problem, which is a numeric computational problem not even known to be in NP and the barrier for polynomial-time tractability for numerous problems in computational geometry, such as the Euclidean shortest path problem, has a polynomial-time reduction to PosSLP [7].
+Mean payoff games are turn-based two-player games on graphs [26, 61]; it is a famous open problem in theoretical computer science whether there is a polynomial-time algorithm to decide for a given graph which of the players has a winning strategy.
+Finding the winning region in mean payoff games is polynomial-time equivalent to various other computational problems in theoretical computer science, for instance scheduling with and-or-constraints [49], the max-atoms problem [14], as well as solvability of max-plus systems and tropical linear feasibility [2, 1].
+The latter can be seen as the tropical analog of testing feasibility of linear inequalities (which for classical geometry is known to be in P, e.g., via the ellipsoid method [42, 37]).
+Furthermore, there is a simple reduction from parity games to mean payoff games by using the priorities of the parity game with a suitably large basis [53].
+Parity games are polynomial-time equivalent to the model-checking problem of the propositional $\mu$-calculus [28], which has been called ‘the most important logics in model checking’ [19].
+For parity games, a quasipolynomial algorithm has been found recently [20].
+Subsequent work indicates that the various quasipolynomial methods that have been obtained lately for parity games [46, 40, 45] do not extend to mean payoff games [22].
+Mean payoff games can be generalised to stochastic mean payoff games (sometimes called $2\frac{1}{2}$-player games), where the graph might contain stochastic nodes.
+If the strategy of one of the two players in such a game is fixed, the game turns into a Markov decision process, for which polynomial-time algorithms based on linear programming are known [54].
+Stochastic mean payoff games are equivalent under polynomial-time Turing reductions to Condon’s simple stochastic games [8] (we may even assume that the simple stochastic games are stopping, see the discussion in Section˜2.2), which are known to be in the intersection of NP and coNP [23].
+There are many other variants of games that all have the same complexity, see [8].
+One approach to analyze mean payoff games consists in reducing these games to constraint satisfaction problems (CSPs).
+More precisely, it is known that mean payoff games are equivalent to the max-atom problem [9, 49], while the more expressive stochastic mean payoff games can be reduced to max-plus-average constraints, which are still in NP $\cap$ coNP [17].
+In the conference version of this article, we introduced a fragment of max-plus-average constraints, called max-average constraints.
+There is a polynomial-time reduction from stopping simple stochastic games to max-average constraints (Section˜3).
+Max-average constraints have the advantage that they can be further reduced to the feasibility problem for semidefinite programs, as we will show here.
+In fact, with some extra effort we can also show this for the entire class of max-plus-average constraints.
+This implies that all the mentioned computational problems from the second cluster can be reduced to semidefinite programming.
+See Figure˜1 for an overview of the mentioned computational problems and their relationship.
 Figure 1: Computational problems that are not known to be in P and not known to be NP-hard, and their relationship.
 Arcs indicate polynomial-time (many-one) reductions, the dotted arc indicates a polynomial-time Turing reduction.
-We introduce a fragment of max-plus-average constraints, which we call max-average constraints.
-There is a polynomial-time reduction from stopping simple stochastic games to max-average constraints (Section 3).
-Max-average constraints have the advantage that they can be further reduced to the feasibility problem for semidefinite programs.
-This implies that all the mentioned computational problems from the second cluster can be reduced to semidefinite programming.
-See Figure 1 for an overview of the mentioned computational problems and their relationship.
-The reduction uses a similar idea as the reduction of Schewe [Sch09] to reduce parity games and mean payoff games to linear programming, which is, however, not a proper polynomial-time reduction.
-The idea is to replace constraints of the form $x_{0}=\max(x_{1},\dots,x_{k})$ by $x_{0}=\log_{b}(b^{x_{1}}+\dots+b^{x_{k}})$ for a sufficiently large $b$.
-Following [AGS18], we extend this idea to max-average constraints and obtain a proper polynomial-time reduction to non-Archimedean SDPs.
+Our reduction consists of two steps.
+The first step uses a similar idea as the reduction of Schewe [57] to reduce parity games and mean payoff games to linear programming, which is, however, not a proper polynomial-time reduction.
+More precisely, the idea of this first step is to replace constraints of the form $x_{0}=\max(x_{1},\dots,x_{k})$, which occur in the max-atom problem, by $x_{0}=\log_{b}(b^{x_{1}}+\dots+b^{x_{k}})$ for a sufficiently large $b$.
+This reduction is not a polynomial-time reduction since the numbers $b^{x_{i}}$ may have exponential bit-sizes for any $x$ that solves the original problem.
+Following [6], we extend this idea to max-plus-average constraints and obtain a proper polynomial-time reduction to non-Archimedean SDPs, which are SDPs defined over a field of formal power series.
 From there, we use bounds from quantifier elimination results to translate non-Archimedean SDPs to real SDPs.
 However, we still have the obstacle that these SDPs involve coefficients of doubly exponential size, which means that they have exponential representation size and do not lead to a polynomial-time reduction.
-We overcome this difficulty by using the duality of semidefinite programming to find small expressions that define coefficients of doubly exponential size, which combined with the above ideas leads to a polynomial-time reduction from max-average constraints to (real) SDPs.
+In the second step of our reduction, we overcome this difficulty by using the duality of semidefinite programming to find small expressions that define coefficients of doubly exponential size, which combined with the above ideas leads to a polynomial-time reduction from max-plus-average constraints to (real) SDPs.
 Our results imply that if there is a polynomial-time algorithm for SDPs, then all the mentioned problems in the second cluster can be solved in polynomial time as well.
 Conversely, our reduction can be used to translate interesting families of instances for (stochastic) mean payoff games or parity games into SDP instances, which might yield interesting families of instances for (algorithmic approaches to) semidefinite programming.
-For linear programming, this idea turned out to be very fruitful: certain instances of parity games have been used to prove exponential lower bounds for several famous pivoting rules for the simplex method in linear programming [FHZ11, Fri11, FHZ14, AF17, DFH23].
+For linear programming, this idea turned out to be very fruitful: certain instances of parity games have been used to prove exponential lower bounds for several famous pivoting rules for the simplex method in linear programming [33, 32, 34, 10, 25].
 However, these results were not based on a general polynomial-time reduction from parity games to linear programming (which is not known!).
@@
 There are multiple works that reduce games to continuous optimization problems.
-As mentioned above, one-player mean payoff games (Markov decision processes) can be reduced to linear programming [Put05], which implies that they can be solved in polynomial time.
-For the case of two-player deterministic mean payoff games, Schewe [Sch09] proposed a reduction to linear programs with coefficients of exponential representation size.
-Such programs can also be interpreted as linear programs over non-Archimedean ordered fields [ABGJ15, ABGJ14, ABGJ21].
-The idea of Schewe was generalized to stochastic games in [BEGM17], where the authors prove that stochastic mean payoff games can be encoded by convex programs with constraints of exponential encoding size.
+As mentioned above, one-player mean payoff games (Markov decision processes) can be reduced to linear programming [54], which implies that they can be solved in polynomial time (but it is open whether there is a strongly polynomial algorithm).
+For the case of two-player deterministic mean payoff games, Schewe [57] proposed a reduction to linear programs with coefficients of exponential representation size.
+Such programs can also be interpreted as linear programs over non-Archimedean ordered fields [4, 3, 5].
+The idea of Schewe was generalized to stochastic games in [18], where the authors prove that stochastic mean payoff games can be encoded by convex programs with constraints of exponential encoding size.
 Such programs can be solved by the ellipsoid method in pseudo-polynomial time provided that the number of stochastic nodes is fixed.
-The non-Archimedean approach was generalized from deterministic games to stochastic games in the work [AGS18], which proves that stochastic mean payoff games can be reduced to semidefinite programs over non-Archimedean fields.
-The paper [AGS18] serves as a basis for our results.
-These reductions express mean payoff games as convex optimization problems, but these problems are “non-standard” in the sense that they either use constraints of exponential encoding size or are expressed over non-Archimedean fields.
+The non-Archimedean approach was generalized from deterministic games to stochastic games in the work [6], which proves that stochastic mean payoff games can be reduced to semidefinite programs over non-Archimedean fields.
+The paper [6] serves as a basis for our results.
+These reductions express mean payoff games as convex optimization problems, but the resulting problems are “non-standard” in the sense that they either use constraints of exponential encoding size or are expressed over non-Archimedean fields.
 There is yet another line of work deriving polynomial-time reductions from mean payoff games to classical optimization problems which are not convex.
-In particular, [Con93] expressed simple stochastic games as non-convex quadratic programs.
-This reduction is further studied in [KRSW22].
-Another reduction to (generalized) complementarity problems with $P$-matrices was proposed in [GR05, SV06] and extended in subsequent works [JS08, FJS10, HIJ13].
-Such linear complementarity problems form a subclass of quadratic programs and belong to the complexity class UniqueEOPL [FGMS20].
+In particular, [24] expressed simple stochastic games as non-convex quadratic programs.
+This reduction is further studied in [43].
+Another reduction to (generalized) complementarity problems with $P$-matrices was proposed in [36, 59] and extended in subsequent works [41, 31, 38].
+Such linear complementarity problems form a subclass of quadratic programs and belong to the complexity class UniqueEOPL [30].
 However, they are in general non-convex and can therefore not be expressed by semidefinite programs.
@@
 The set of natural numbers $\{0,1,2,\dots\}$ is denoted by $\mathbb{N}$}, and the set of real numbers by $\mathbb{R}$}.
-We assume familiarity with first-order logic; see, e.g., [Hod97].
+We assume familiarity with first-order logic; see, e.g., [39].
 Let $\tau$ be a signature, i.e., a set of function and relation symbols.
-A first-order $\tau$-formula is a formula build from the Boolean connectives, atomic $\tau$-formulas and quantification in the usual way.
+A first-order $\tau$-formula is a formula built from the Boolean connectives, atomic $\tau$-formulas and quantification in the usual way.
 If $A$ is a $\tau$-structure and $\phi$ is a $\tau$-formula with free variables $x_{1},\dots,x_{k}$, then $\phi$ defines over $A$ the relation
 $\{(a_{1},\dots,a_{k})\in A^{k}\mid A\models\phi(a_{1},\dots,a_{k})\}.$
-A set if called semialgebraic if it has a first-order definition over the field $({\mathbb{R}};+,\cdot)$, allowing real parameters.
+A set is called semialgebraic if it has a first-order definition over the field $({\mathbb{R}};+,\cdot)$, allowing real parameters (we do not need the order in the signature since it is definable from addition and multiplication).
 A first-order formula $\phi(x_{1},\dots,x_{k})$ with free variables $x_{1},\dots,x_{k}$ is called primitive positive if it is of the form
@@
 2.1 Semidefinite Programming
-For a symmetric matrix $S\in{\mathbb{R}}^{m\times m}$ we write $S\succeq 0$ if $S$ is positive semidefinite, i.e., if $x^{\top}Sx\geq 0$ for all $x\in{\mathbb{R}}^{m}$, and $S\succ 0$ if it is positive definite, i.e., if $x^{\top}Sx>0$ for all $x\in{\mathbb{R}}^{m}\setminus\{0\}$.
+For a matrix $S\in{\mathbb{R}}^{m\times m}$ we write
+  • $S\succeq 0$ if $S$ is positive semidefinite, i.e., $S$ is symmetric and $x^{\top}Sx\geq 0$ for all $x\in{\mathbb{R}}^{m}$, and
+  • $S\succ 0$ if it is positive definite, i.e., $S$ is symmetric and $x^{\top}Sx>0$ for all $x\in{\mathbb{R}}^{m}\setminus\{0\}$.
 A semidefinite program (SDP) is an optimisation problem of the following form:
@@
 where $A_{1},\dots,A_{n},B\in{\mathbb{R}}^{m\times m}$ are symmetric and $c_{1},\dots,c_{n}\in{\mathbb{R}}$.
-If the data $A_{1},\dots,A_{n},B,c_{1},\dots,c_{n}$ is rational, then the problem of deciding whether the supremum exists, and if so, to compute the supremum, is a computational problem known as the semidefinite programming problem.
+If the data $A_{1},\dots,A_{n},B$, $c_{1},\dots,c_{n}$ is rational, then the problem of deciding whether the infimum exists, and if so, to compute the infimum, is a computational problem known as the semidefinite programming problem.
 The representation size of the SDP then consists of the sum of the bit lengths of all the rational numbers in the input.
-The expression $\sum_{i=1}^{n}x_{i}A_{i}-B\succeq 0$ in Section 2.1 is called a linear matrix inequality (LMI), and defines the feasible region of the SDP, which is called a spectrahedron.
+The expression $\sum_{i=1}^{n}x_{i}A_{i}-B\succeq 0$ in Section˜2.1 is called a linear matrix inequality (LMI), and defines the feasible region of the SDP, which is called a spectrahedron.
 The representation size of LMIs is defined analogously as for SDPs.
@@
 Given symmetric matrices $A_{1},\dots,A_{n},B\in{\mathbb{Q}}^{m\times m}$ the feasibility problem for semidefinite programs asks to decide whether there is a solution $x\in\mathbb{R}^{n}$ to an LMI
-$x_{1}A_{1}+\dots+x_{n}A_{n}-B\succeq 0\enspace.$
+$x_{1}A_{1}+\dots+x_{n}A_{n}-B\succeq 0\kern 5.0pt.$
 Note that spectrahedra are convex and semialgebraic.
-It has been conjectured that every convex semialgebraic set is the projection of a spectrahedron (also called a spectrahedral shadow), but this turned out to be false [Sch18] (see [BKT24] for an alternative proof of an even stronger inexpressibility result).
+It has been conjectured that every convex semialgebraic set is the projection of a spectrahedron (also called a spectrahedral shadow), but this turned out to be false [56] (see [15] for an alternative proof of an even stronger inexpressibility result).
 It is easy to see that the set of spectrahedral shadows is closed under primitive positive definability, and that for every primitive positive formula $\phi$ we may find an LMI for the spectrahedral shadow defined by $\phi$ whose representation size is bounded by the sum of the representation sizes of conjuncts of $\phi$.
 We also note that the definition of a semidefinite program makes sense over any real closed field, not only over the real numbers.
-In particular, it is meaningful to consider semidefinite programs over non-Archimedean real closed fields, such as the field of Puiseux series which we discuss in Section 4.
+In particular, it is meaningful to consider semidefinite programs over non-Archimedean real closed fields, such as the field of Puiseux series which we discuss in Section˜4.
 The (Frobenius) inner product of two matrices $S,T\in{\mathbb{R}}^{m\times m}$ is defined as $\langle S,T\rangle:=\text{trace}(ST)=\sum_{i,j}S_{ij}T_{ij}$.
-The dual program to the SDP from Section 2.1 (to which we refer as the primal) is given by
+The dual program to the SDP from Section˜2.1 (to which we refer as the primal) is given by
 $\sup_{Y\succeq 0}$ $\langle B,Y\rangle$ (D)
@@
 Moreover, both primal and dual might not have optimal solutions, i.e., the supremum may not be a maximum and the infimum might not be a minimum.
-See [Ram97] for a more complicated, but exact dual.
+See [55] for a more complicated, but exact dual.
 However, for us the following sufficient condition for a vanishing duality gap will be enough.
-Theorem 2.2 (see Theorem 2.4.1 in [BTN01]).
+Theorem 2.2 (see Theorem 2.4.1 in [12] and the explanations on page 142 and 143).
 If the primal SDP is strictly feasible and bounded from below, then the dual program attains its supremum and there is no duality gap.
 2.2 Stochastic Games
-A stochastic mean payoff game is a two-player game played on a finite edge-weighted directed graph $\vec{\mathcal{G}}=(V,E,r)$, i.e., $V$ is a finite set of nodes, $E\subseteq V^{2}$ is a set of edges, and $r\colon E\to{\mathbb{Q}}$ is a weight function.
+A stochastic mean payoff game is a two-player game played on a finite edge-weighted directed graph $\vec{\mathcal{G}}=(V,E,r,p)$, i.e., $V$ is a finite set of nodes, $E\subseteq V^{2}$ is a set of edges, $r\colon E\to{\mathbb{Q}}$ is a weight function, and $p\colon E\to{\mathbb{Q}}\cap[0,1]$ specifies edge probabilities.
 The nodes of the graph are partitioned into three subsets, $V=V_{\mathrm{Min}}\sqcup V_{\mathrm{Max}}\sqcup V_{\mathrm{Rand}}$.
-The nodes in $V_{\mathrm{Min}}$ are controlled by player Min, the nodes in $V_{\mathrm{Max}}$ are controlled by player Max, and the nodes in $V_{\mathrm{Rand}}$ are controlled by nature.
+The nodes in $V_{\mathrm{Min}}$ are controlled by player Min, the nodes in $V_{\mathrm{Max}}$ are controlled by player Max, and the nodes in $V_{\mathrm{Rand}}$ are controlled by nature; the probabilities of the outgoing edges of each vertex in $V_{\mathrm{Rand}}$ sum up to one.
 The players play the game by moving a pawn on the graph.
@@
 Analogously, player Max chooses the next move when the pawn is on a node $v\in V_{\mathrm{Max}}$.
-When the pawn is on a node $v\in V_{\mathrm{Rand}}$, the next move is chosen by nature randomly, according to some probability distribution on the outgoing edges of $v$.
-This probability distribution is known to both players and does not change during the game.
+When the pawn is on a node $v\in V_{\mathrm{Rand}}$, the next move is chosen by nature randomly according to the probability distribution given by $p$.
 To ensure that players can always make a move, we suppose that all nodes of the graph have at least one outgoing edge.
@@
 The objective of player Max is to maximize the mean payoff reward while the objective of player Min is to minimize this reward.
-It is known that both players in a stochastic mean payoff game have optimal policies [LL69].
+It is known that both players in a stochastic mean payoff game have optimal policies [47].
 In other words, there exist policies $(\sigma^{*},\tau^{*})$ and a vector $\chi\in\mathbb{R}^{V}$ such that for every $u$ and every $(\sigma,\tau)$ we have the inequality
@@
 These states are absorbing, i.e., $\mathrm{Win}$ has only one outgoing edge that goes back to $\mathrm{Win}$ and $\mathrm{Lose}$ has only one outgoing edge that goes back to $\mathrm{Lose}$.
-The weight of the edge going from $\mathrm{Win}$ to $\mathrm{Win}$ is equal to $1$, while the weight of all of the other edges of the graph is equal to 00.
+The weight of the edge going from $\mathrm{Win}$ to $\mathrm{Win}$ is equal to $1$, while the weight of all of the other edges of the graph is equal to 0.
 In this case, the mean payoff reward of player Max is simply the probability that the pawn reaches the winning state, $g(\sigma,\tau)_{u}=\mathbb{P}^{u}_{\sigma,\tau}(\text{Reach}(\mathrm{Win}))$.
@@
 The computational problem associated with stochastic mean payoff games is the following: given a stochastic mean payoff game, find its value and a pair of optimal policies.
-In [AM09], Andersson and Miltersen proved that this problem is polynomial-time (Turing) reducible to the problem of finding the value of a stopping simple stochastic game.
+In [8], Andersson and Miltersen proved that this problem is polynomial-time (Turing) reducible to the problem of finding the value of a stopping simple stochastic game.
 By a binary search argument, the problem of computing the value of a stopping simple stochastic game can be further reduced to the decision problem, in which we want to decide whether the value of a given state of such a game is at least $1/2$.
-The details of this reduction can be found, e.g., in [AGS18] or [Sko18, Section 7.1].
-Theorem 2.3 ([AGS18, Lemma 48]).
+The details of this reduction can be found, e.g., in [6] or [58, Section 7.1].
+Theorem 2.3 ([6, Lemma 48]).
 The problem of computing the value and a pair of optimal policies of a stochastic mean payoff game is polynomial-time (Turing) reducible to the problem of deciding whether a given state of a stopping simple stochastic game has value at least $1/2$.
 3 From stochastic games to max-average constraints
-In this section, we focus on the decision problem stated in Theorem 2.3 above: given a stopping simple stochastic game we want to decide whether the value of a given state is at least $1/2$.
-This problem can be reduced to max-average constraint satisfaction problems.
+In this section, we focus on the decision problem stated in Theorem˜2.3 above: given a stopping simple stochastic game we want to decide whether the value of a given state is at least $1/2$.
+This problem can be reduced to max-average constraint satisfaction problems (max-average CSPs).
 The ideas for this reduction are not new, but we are not aware of an explicit reference for this fact, and we therefore present a short proof in this section.
-In the next section, we will show that max-average CSPs then further reduce to the feasibility problem of non-Archimedean semidefinite programs.
-The fact that stochastic games can be encoded as non-Archimedean SDPs was proven in [AGS18] and our reduction is very similar.
+In the next section, we will show that max-average CSPs, and even the larger class of max-plus-average CSPs, then further reduce to the feasibility problem of non-Archimedean semidefinite programs.
+The fact that stochastic games can be encoded as non-Archimedean SDPs was proven in [6] and our reduction is very similar.
 While their approach relies on reducing stochastic games to ergodic games, we omit this step and obtain a slightly more straightforward reduction.
 Definition 3.1.
-An instance of a max-average constraint satisfaction problem is a conjunction of constraints, each of which has one of the following forms:
+An instance of a max-plus-average CSP is a conjunction of constraints, each of which has one of the following forms:
   1. 1.
@@
   2. 2.
+    $x_{0}\leq c+x_{1}$ where $c$ is a rational constant,
+  3. 3.
     $x_{0}\leq\frac{x_{1}+x_{2}}{2}$,
-  3. 3.
+  4. 4.
     $x_{0}=c$ where $c$ is a rational constant.
 The computational task is to determine whether such a conjunction has a solution over $\mathbb{Q}\cup\{-\infty\}$ (equivalently, over $\mathbb{R}\cup\{-\infty\}$).
-Note that max-average CSP instances may also use constraints of the form $x_{0}\leq\min(x_{1},\dots,x_{k})$ where $k\in\mathbb{N}$, $x_{0}\leq c$, or $x_{0}\geq c$ for some rational constant $c$, because the respective relations are primitively positively definable over the other relations.
+If the instance does not contain constraints of the form $2.$, we call it a max-average instance.
+Note that max-average CSP instances may also use constraints of the form $x_{0}\leq\min(x_{1},\dots,x_{k})$ where $k\in\mathbb{N}$, $x_{0}\leq c$, or $x_{0}\geq c$ for some rational constant $c$, because the respective relations are primitively positively definable using max-average constraints.
 We now show that the problem of solving stochastic mean payoff games can be reduced to max-average CSPs.
-A reduction to slightly more general CSPs, including constraints of the form $x_{0}\leq c+x_{1}$, follows from [AGS18, Theorem 18] and is based on a Collatz–Wielandt property of order-preserving and additively homogeneous maps.
-However, for our purposes, it is useful that the reduction does not use these additional constraints.
-Instead, we give a proposition that characterizes the value of a stopping simple stochastic game.
-It was already proven in [Con92] that this value is a unique fixed point of an operator involving min, max, and averaging operations.
+A reduction to max-plus-average CSPs follows from [6, Theorem 18] and is based on a Collatz–Wielandt property of order-preserving and additively homogeneous maps.
+We provide a more direct reduction to the smaller class of CSPs without constraints of the form $x_{0}\leq c+x_{1}$ as an intermediate step for our reduction.
+To do so, we give a proposition that characterizes the value of a stopping simple stochastic game.
+It was already proven in [23] that this value is a unique fixed point of an operator involving min, max, and averaging operations.
 By replacing the fixed point equations with inequalities, we obtain a set of certificates which prove that the value is at least $1/2$.
-This observation is proven, e.g., in [BBDFS23] and gives inequalities very similar to the inequalities on Shapley operators used in [AGS18].
+This observation is proven, e.g., in [13] and gives inequalities very similar to the inequalities on Shapley operators used in [6].
 Proposition 3.2.
@@
 Then, the value of a state $k\in[n]$ is at least $1/2$ if and only if the following system of inequalities (an instance of a max-average constraint satisfaction problem) has a solution $x\in\mathbb{Q}_{\geq 0}^{n}$:
-$x_{i}$ $\leq\max_{(i,j)\in E}\{x_{j}\}$ for every $\text{for every }\in V_{\mathrm{Max}},$ (1)
-$x_{i}$ $\leq\min_{(i,j)\in E}\{x_{j}\}$ for every $\text{for every }i\in V_{\mathrm{Min}},$
-$x_{i}$ $\leq\frac{1}{2}\sum_{(i,j)\in E}x_{j}$ for every $\text{for every }i\in V_{\mathrm{Rand}},$
+$x_{i}$ $\leq\max_{(i,j)\in E}\{x_{j}\}$ $\text{for every }i\in V_{\mathrm{Max}},$ (1)
+$x_{i}$ $\leq\min_{(i,j)\in E}\{x_{j}\}$ $\text{for every }i\in V_{\mathrm{Min}},$
+$x_{i}$ $\leq\frac{1}{2}\sum_{(i,j)\in E}x_{j}$ $\text{for every }i\in V_{\mathrm{Rand}},$
 $x_{\mathrm{Lose}}$ $=0,$
@@
 $x_{k}$ $\geq 1/2.$
-Proof.
-Let $(Q)$ denote the system of inequalities obtained from Eq.
-1 by removing the inequality $x_{k}\geq 1/2$.
-By [Con92] or [BBDFS23, Theorem 77], the value of the game satisfies $(Q)$ with inequalities replaced by equalities.
-Hence, if the value at state $k$ is at least $1/2$, then the value of the game satisfies Eq.
-1.
-Moreover, the value of the game is a rational vector by [Con92, Lemma 2] or [BBDFS23, Corollary 18].
-Conversely, if a point $x\in\mathbb{Q}_{\geq 0}^{n}$ satisfies $(Q)$, then [BBDFS23, Lemma 83] implies that for all $i\in[n]$, the entry $x_{i}$ is not greater than the value of the game at state $i$.
+Proof 3.3.
+Let $(Q)$ denote the system of inequalities obtained from Equation˜1 by removing the inequality $x_{k}\geq 1/2$.
+By [23] or [13, Theorem 77], the value of the game satisfies $(Q)$ with inequalities replaced by equalities.
+Hence, if the value at state $k$ is at least $1/2$, then the value of the game satisfies Equation˜1.
+Moreover, the value of the game is a rational vector by [23, Lemma 2] or [13, Corollary 18].
+Conversely, if a point $x\in\mathbb{Q}_{\geq 0}^{n}$ satisfies $(Q)$, then [13, Lemma 83] implies that for all $i\in[n]$, the entry $x_{i}$ is not greater than the value of the game at state $i$.
 Hence, if we further suppose that $x_{k}\geq 1/2$, then the value of the game at state $k$ is at least $1/2$.
-∎
-4 From max-average constraints to non-Archimedean SDPs
-By “lifting” the inequalities Eq.
-1 to Puiseux series, we get the reduction from max-average constraint satisfaction problems to non-Archimedean SDPs.
+4 From max-plus-average constraints to non-Archimedean SDPs
+By “lifting” the inequalities in Definition 3.1 to Puiseux series, we get the reduction from max-plus-average constraint satisfaction problems to non-Archimedean SDPs.
 A (formal real) Puiseux series is a series of the form
@@
 Moreover, they can be ordered by defining $\bm{x}\geq 0$ if $c_{0}\geq 0$ and $\bm{x}\geq\bm{y}$ if $\bm{x}-\bm{y}\geq 0$.
-It is known that Puiseux series form a real closed field, see, e.g., [BPR06, Section 2.6] for a detailed discussion.
+It is known that Puiseux series form a real closed field, see, e.g., [11, Section 2.6] for a detailed discussion.
 [^1] In this work, we use a convention in which one interprets $t$ as a “very large” parameter.
 Many authors use the opposite but equivalent convention in which $t$ is a very small parameter.
-This explains the differences in notation between our work and [BPR06].
+This explains the differences in notation between our work and [11].
+This field is non-Archimedean, meaning that the order does not satisfy the Archimedean property: the series $\bm{x}(t)=t$ is larger than any natural number, i.e., $t\geq m$ for all $m\in\mathbb{N}$.
 We denote the field of Puiseux series by $\mathbb{K}$.
@@
 Theorem 4.1.
-An instance $\phi$ with variables $x_{1},\dots,x_{n}$ of a max-average constraint satisfaction problem has a solution in ({}$(\mathbb{Q}\cup\{-\infty\})^{n}$ if and only if the following system of inequalities $S_{\phi}(t)$, which is a (quadratic) semidefinite program over Puiseux series, has a solution in $\mathbb{K}^{n}$: for every variable $x_{i}$ of $\phi$ we have a variable $\bm{x}_{i}$, and the inequalities are
+An instance $\phi$ with variables $x_{1},\dots,x_{n}$ of a max-plus-average constraint satisfaction problem has a solution in $(\mathbb{Q}\cup\{-\infty\})^{n}$ if and only if the following system of inequalities $S_{\phi}(t)$, which is a (quadratic) semidefinite program over Puiseux series, has a solution in $\mathbb{K}^{n}$: for every variable $x_{i}$ of $\phi$ we have a variable $\bm{x}_{i}$, and the inequalities are
 $\bm{x}_{i_{0}}$ $\leq\bm{x}_{i_{1}}+\cdots+\bm{x}_{i_{k}}$ for every constraint $x_{i_{0}}\leq\max(x_{i_{1}},\dots,x_{i_{k}})$ in $\phi$, (3)
+$\bm{x}_{i_{0}}$ $\leq t^{c}\bm{x}_{i_{1}}$ for every constraint $x_{i_{0}}\leq c+x_{i_{1}}$,
 $\bm{x}^{2}_{i_{0}}$ $\leq\bm{x}_{i_{1}}\bm{x}_{i_{2}}$ for every constraint $x_{i_{0}}\leq\frac{x_{i_{1}}+x_{i_{2}}}{2}$ in $\phi$,
@@
 $\bm{x}_{i_{0}}$ $\geq 0$ for every $i_{0}\in\{1,\dots,n\}$.
-Proof.
-Suppose that $\bm{x}=(\bm{x}_{1},\dots,\bm{x}_{n})\in\mathbb{K}^{n}$ satisfies Eq.
-3.
-Then, $\operatorname*{\mathsf{val}}(\bm{x})$ belongs to ({}$(\mathbb{Q}\cup\{-\infty\})^{n}$ and satisfies $\phi$ by the properties of the valuation function Eq.
-2.
-Conversely, if $(x_{1},\dots,x_{n})\in(\mathbb{Q}\cup\{-\infty\})^{n}$ satisfies $\phi$, then $(t^{x_{1}},\dots,t^{x_{n}})$ (with the convention that $t^{-\infty}=0$) satisfies Eq.
-3.
+Proof 4.2.
+Suppose that $\bm{x}=(\bm{x}_{1},\dots,\bm{x}_{n})\in\mathbb{K}^{n}$ satisfies Equation˜3.
+Then, $\operatorname*{\mathsf{val}}(\bm{x})$ belongs to $(\mathbb{Q}\cup\{-\infty\})^{n}$ and satisfies $\phi$ by the properties of the valuation function Equation˜2.
+Conversely, if $(x_{1},\dots,x_{n})\in(\mathbb{Q}\cup\{-\infty\})^{n}$ satisfies $\phi$, then $(t^{x_{1}},\dots,t^{x_{n}})$ (with the convention that $t^{-\infty}=0$) satisfies Equation˜3.
 This amounts to the feasibility problem for a semidefinite program: the linear (in)equalities can be expressed by diagonal matrices and each quadratic condition in combination with the non-negativity of all variables is just the positive semidefiniteness of a symmetric 2-by-2 matrix.
-Conjunctions of such constraints can be expressed by LMIs for block diagonal matrices.
-∎
+Indeed, the symmetric matrix $\begin{pmatrix}x_{1}&x_{3}\\ x_{3}&x_{2}\end{pmatrix}$ is positive semidefinite if and only if $x_{1}\geq 0$, $x_{2}\geq 0$, and $x_{1}x_{2}\geq x_{3}^{2}$.
+(More generally, a symmetric matrix is positive semidefinite if and only if all of its principal minors are nonnegative.) These are precisely the inequalities that appear in Equation˜3, together with some linear inequalities.
+A conjunction of such inequalities can then be expressed by a single LMI given by block-diagonal matrices (a symmetric block-diagonal matrix is positive semidefinite if and only if every block is positive semidefinite).
 5 From non-Archimedean to real SDPs
-To conclude, we aim to reduce the non-Archimedean semidefinite program from Eq.
-3 to a real semidefinite program.
+To conclude, we aim to reduce the non-Archimedean semidefinite program from Equation˜3 to a real semidefinite program.
 We will do this in two steps.
-First, we replace the formal parameter $t$ by a real number that is sufficiently large (doubly exponential in the size of the CSP).
+First, we replace the formal parameter $t$ by a real number that is sufficiently large (doubly exponential in the size of the instance).
 In fact, any non-Archimedean semialgebraic feasiblity problem can be reduced to a semialgebraic feasibility problem over the reals by replacing $t$ with a sufficiently large real number.
 This follows from the quantifier elimination in real closed fields coupled with the definition of the order in Puiseux series.
-By using the effective bounds on quantifier elimination discussed, e.g., in [BPR06], we can give a bound on this “sufficiently large” number.
+By using the effective bounds on quantifier elimination discussed, e.g., in [11], we can give a bound on this “sufficiently large” number.
 Second, we will show that this SDP with very large coefficients can be reduced to another SDP that has coefficients of polynomial bit-size.
-Before presenting the proofs, we note that a reduction from non-Archimedean SDPs coming from games to real SDPs with large coefficients was already studied in [AGS18, Theorem 36].
-However, the bounds given in [AGS18] rely on an additional assumption on the underlying game.
-In particular, they cannot be applied directly to the SDP given in Eq.
-3.
+Before presenting the proofs, we note that a reduction from non-Archimedean SDPs coming from games to real SDPs with large coefficients was already studied in [6, Theorem 36].
+However, the bounds given in [6] rely on an additional assumption on the underlying game.
+In particular, they cannot be applied directly to the SDP given in Equation˜3.
 For max-average systems coming from games, one could overcome this restriction by perturbing the underlying game slightly.
 Nevertheless, the bounds coming from quantifier elimination which we present below do not require any additional assumptions and can be of independent interest.
+We recall that Tarski’s quantifier elimination allows to eliminate quantifiers of first-order $\tau$-formulas, with signature $\tau$ given by $\tau=\{+,\cdot,0,1,\leq\}$.
 Theorem 5.1.
-Let $P_{1},\dots,P_{m}\in\mathbb{Z}[x_{0},x_{1},\dots,x_{n}]$ be a collection of polynomials with integer coefficients and let {,,<,>,=,}$\diamond\in\{\leq,\geq,<,>,=,\neq\}^{m}$ be a sign pattern.
+Let $P_{1},\dots,P_{m}\in\mathbb{Z}[x_{0},x_{1},\dots,x_{n}]$ be a collection of polynomials with integer coefficients and let $\diamond\in\{\leq,\geq,<,>,=,\neq\}^{m}$ be a sign pattern.
 Consider the semialgebraic set over Puiseux series defined by
@@
 More precisely, there exists an absolute constant $C>0$ with the following property.
-If the maximal degree of any polynomial $P_{i}$ is bounded by $d\geq 2$ and $s$ denotes the maximal bit-size of any coefficient of $P_{i}$, then for $L>2^{sd^{Cn}}$ either all $S(L)$ are empty or all $S(L)$ are nonempty.
+If the maximal degree of any polynomial $P_{i}$ is bounded by $d\geq 2$ and $s$ denotes the maximal bit-size of any coefficient of $P_{i}$, then either for all $L>2^{sd^{Cn}}$ we have $S(L)=\emptyset$, or for all $L>2^{sd^{Cn}}$ we have $S(L)\neq\emptyset$.
 In the former case $\bm{S}$ is empty and in the latter case $\bm{S}$ is nonempty.
-Proof.
-Consider the first-order formula $\phi(x_{0})$ in the language of ordered rings given by
+Proof 5.2.
+Both sets $S(L)$ and $\bm{S}$ can be defined by the same first-order formula with a free variable $x_{0}$ that can be substituted by a real number $L$ (to obtain $S(L)$) or a Puiseux series $t$ (to obtain $\bm{S}$).
+Let $\tau=\{+,\cdot,0,1,\leq\}$ and consider the first-order $\tau$-formula $\phi(x_{0})$ given by
 $\exists x_{1},\dots,\exists x_{n}\operatorname*{\bigwedge}_{i=1}^{m}(P_{i}(x_{0},x)\diamond_{i}0)\,.$
-By Tarski’s quantifier elimination theorem [BPR06, Theorem 2.77], there exists a quantifier-free formula $\psi(x_{0})$ that is equivalent to $\phi(x_{0})$ in the theory of real closed fields.
-The bounds on quantifier elimination given in [BPR06, Theorem 14.6] show that $\psi(x_{0})$ can be taken to be of the form
+By Tarski’s quantifier elimination theorem [11, Theorem 2.77], there exists a quantifier-free formula $\psi(x_{0})$ that is equivalent to $\phi(x_{0})$ in the theory of real closed fields.
+The bounds on quantifier elimination given in [11, Theorem 14.6] show that $\psi(x_{0})$ can be taken to be of the form
 $\operatorname*{\bigvee}_{i=1}^{I}\operatorname*{\bigwedge}_{j=1}^{J_{i}}\left(\operatorname*{\bigvee}_{k=1}^{N_{ij}}\operatorname*{\mathrm{sign}}(Q_{ijk}(x_{0}))=\delta_{ijk}\right)\,$,
@@
 Furthermore, every $Q_{ijk}$ has degree bounded by $d^{O(n)}$ and every coefficient of every $Q_{ijk}$ has bit-size bounded by $sd^{O(n)}$.
-Hence, the Cauchy bound [BPR06, Corollary 10.4] implies that every real root of $Q_{ijk}$ has absolute value bounded by $2^{sd^{O(n)}}$ (provided that $Q_{ijk}$ is a non-zero polynomial).
+Hence, the Cauchy bound [11, Corollary 10.4] implies that every real root of $Q_{ijk}$ has absolute value bounded by $2^{sd^{O(n)}}$ (provided that $Q_{ijk}$ is a non-zero polynomial).
 In particular, the sign of $Q_{ijk}(L)$ is the same for all $L>2^{sd^{O(n)}}$.
-Therefore, either $S(L)$ are empty for all such $L$ or they are nonempty for all such $L$.
+Therefore, either $S(L)$ is empty for all such $L$ or $S(L)$ is nonempty for all such $L$.
 Moreover, $Q_{ijk}(t)\in\mathbb{K}$ is a Puiseux series with finitely many terms.
@@
 Hence, $\bm{S}$ is nonempty if and only if $S(L)$ is nonempty.
-∎
 In the next proposition and later, we use the following notation.
@@
 Proposition 5.3.
-There exists an absolute integer constant $M>0$ such that an instance $\phi$ of a max-average constraint satisfaction problem has a solution if and only if the real SDP $S_{\phi}(L)$ obtained from $S_{\phi}(t)$ (see Theorem 4.1) by replacing $t$ with $L>2^{2^{Mn\log_{2}W}}$, is feasible.
-Proof.
-Consider the system of inequalities given in Eq.
-3.
-Since $\bm{x}_{i}\geq 0$ for all $i$, any equality of the form $\bm{x}_{i}=t^{c_{j}}=t^{p_{j}/q_{j}}$ can be replaced with an equivalent equality $\bm{x}_{i}^{q_{j}}=t^{p_{j}}$ if $p_{j}\geq 0$ and $t^{-p_{j}}\bm{x}_{i}^{q_{j}}=1$ if $p_{j}<0$.
-If we further replace the parameter $t$ by a variable symbol $x_{0}$, Eq.
-3 becomes a system of polynomial inequalities of degree at most $W$ and such that the coefficients of the polynomials appearing in these inequalities are integers in $\{-1,0,1\}$.
-Hence, by Theorem 5.1, there exists $M>0$ such that the system Eq.
-3 has a solution over Puiseux series if and only if the system obtained by replacing $t$ with $L>2^{2^{Mn\log_{2}W}}$ has a solution over the real numbers.
-Therefore, the claim follows from Theorem 4.1.
-∎
-In Proposition 5.2, we obtain a (real) semidefinite program, whose size, however, is exponential in the size of $\phi$ and therefore does not provide a polynomial-time reduction to the semidefinite program feasiblity problem.
-However, by the following result we can also reduce to a semidefinite program of polynomial representation size.
+There exists an absolute integer constant $M>0$ such that an instance $\phi$ of a max-plus-average constraint satisfaction problem has a solution if and only if the real SDP $S_{\phi}(L)$ obtained from $S_{\phi}(t)$ (see Theorem˜4.1) by replacing $t$ with $L>2^{2^{Mn\log_{2}W}}$, is feasible.
+Proof 5.4.
+Consider the system of inequalities given in Equation˜3.
+Since $\bm{x}_{i}\geq 0$ for all $i\in\{1,\dots,n\}$, any equality of the form $\bm{x}_{i}=t^{c_{j}}=t^{p_{j}/q_{j}}$, for an integer $q_{j}\geq 1$, can be replaced with an equivalent equality $\bm{x}_{i}^{q_{j}}=t^{p_{j}}$ if $p_{j}\geq 0$ and $t^{-p_{j}}\bm{x}_{i}^{q_{j}}=1$ if $p_{j}<0$.
+Likewise, any inequality of the form $\bm{x}_{i}\leq t^{p_{j}/q_{j}}\bm{x}_{l}$ can be replaced by the equivalent inequality $\bm{x}_{i}^{q_{j}}\leq t^{p_{j}}\bm{x}_{l}^{q_{j}}$ if $p_{j}\geq 0$ and $t^{-p_{j}}\bm{x}_{i}^{q_{j}}\leq\bm{x}_{l}^{q_{j}}$ if $p_{j}<0$.
+If we further replace the parameter $t$ by a variable symbol $x_{0}$, Equation˜3 becomes a system of polynomial inequalities of degree at most $W$ and such that the coefficients of the polynomials appearing in these inequalities are integers in $\{-1,0,1\}$.
+Hence, by Theorem˜5.1, there exists $M>0$ such that the system Equation˜3 has a solution over Puiseux series if and only if the system obtained by replacing $t$ with $L>2^{2^{Mn\log_{2}W}}$ has a solution over the real numbers.
+Therefore, the claim follows from Theorem˜4.1.
+In Proposition˜5.3, we obtain a (real) semidefinite program, whose size, however, is exponential in the size of $\phi$ and therefore does not provide a polynomial-time reduction to the semidefinite program feasiblity problem.
+We now show how to reduce to a semidefinite program of polynomial representation size.
+This is based on the following result.
 Theorem 5.5.
 For every $n\in{\mathbb{Z}}$, the singleton set $\{2^{n}\}$ is the projection of a spectrahedron whose representation size is polynomial in the bit-size of $n$.
-Proof.
-To prove the claim, we extend the constructions of SDPs with solutions of doubly-exponential size given in [PK97, Ram97].
+Proof 5.6.
+To prove the claim, we extend the constructions of SDPs with solutions of doubly-exponential size given in [52, 55].
 For $n=0$, the set $\{x\in\mathbb{R}\mid x=1\}$ is the desired spectrahedron.
@@
 Then, the dual is of the form $\sup\,\langle B,Y\rangle$ such that $\langle A_{i},Y\rangle=c_{i}$ for every $i\in\{0,\dots,k\}$ and $Y\succeq 0$, where $(c_{0},\dots,c_{k})=(0,\dots,0,1)$.
-Clearly, the primal is strictly feasible and bounded by $2^{n}$ from below, so we may use strong duality (Theorem 2.2) and obtain that $\sup\,\langle B,Y\rangle=2^{n}$ and that the supremum is attained for some $Y$.
+Clearly, the primal is strictly feasible and bounded by $2^{n}$ from below, so we may use strong duality (Theorem˜2.2) and obtain that $\sup\,\langle B,Y\rangle=2^{n}$ and that the supremum is attained for some $Y$.
 Then the following primitive positive formula over spectrahedral sets with one free variable $x_{k}$ defines the singleton relation $\{2^{n}\}$:
-$\exists x_{0},\dots,x_{k-1},y_{1,1},\dots,y_{m,m}:A_{0}x_{0}+\cdots+A_{k}-B\succeq 0$
+$\exists x_{0},\dots,x_{k-1},y_{1,1},\dots,y_{m,m}:A_{0}x_{0}+\cdots+A_{k}x_{k}-B\succeq 0$ (4)
 $\wedge\,\langle A_{0},Y\rangle=0\,\wedge\cdots\wedge\,\langle A_{k-1},Y\rangle=0\,\wedge\,\langle A_{k},Y\rangle=1$ $\wedge Y\succeq 0$
@@
 where $Y$ denotes $(y_{i,j})_{i,j\in\{1,\dots,m\}}$ for $m=2(k+1)$.
-As we have mentioned before, spectrahedral shadows are closed under primitive positive definability, and the representation size of the resulting LMI is bounded by the representation sizes for the conjuncts.
+As we have mentioned before, spectrahedral shadows are closed under primitive positive definability, and the representation size of the resulting LMI is polynomially bounded by the representation sizes for the conjuncts.
+Indeed, the matrices $A_{i}$ and $B$ have coefficients in $0,\pm 1/2,\pm 1,\pm 2$} and are of size $k\times k$ which is polynomial in the bit-size of $n$.
+Hence, the system of matrix inequalities Equation˜4 can be encoded using a number of bits that is polynomial in the bit-size of $n$.
+This system of inequalities can be then encoded by a single block-diagonal LMI as in the proof of Theorem˜4.1.
+This increases the number of necessary bits by a polynomial factor.
 We thus obtain an LMI of representation size that is polynomial in $k$.
-∎
 Corollary 5.7.
 There is a polynomial-time reduction from the max-average constraint satisfaction problem to the feasibility problem for semidefinite programs.
-Proof.
-We have already specified a first reduction to a (real) SDP in Proposition 5.2.
+Proof 5.8.
+We have already specified a first reduction to a (real) SDP in Proposition˜5.3.
 To obtain a proper polynomial-time reduction, we modify this SDP as follows.
-Fix an instance $\phi$; let $c_{i}$, $p_{i}$, $q_{i}$, and $W$ be as introduced before Proposition 5.2.
-Define $D\coloneqq\prod_{i}q_{i}$ and $K\coloneqq 2DW^{Mn}=D2^{1+Mn\log_{2}W}$, where $M$ is the integer constant from Proposition 5.2.
+Fix a max-average instance $\phi$; let $c_{i}$, $p_{i}$, $q_{i}$, and $W$ be as introduced before Proposition˜5.3.
+Define $D\coloneqq\prod_{i}q_{i}$ and $K\coloneqq 2DW^{Mn}=D2^{1+Mn\log_{2}W}$, where $M$ is the integer constant from Proposition˜5.3.
 Furthermore, let $L\coloneqq 2^{K}$.
-By Proposition 5.2, $\phi$ has a solution if and only if $S_{\phi}(L)$ has a solution over the real numbers.
+By Proposition˜5.3, the instance $\phi$ has a solution if and only if $S_{\phi}(L)$ has a solution over the real numbers.
 This SDP has constraints of the form $x_{i}=L^{c_{j}}=2^{c_{j}K}$.
 Since $K$ is an integer divisible by $D$, the number $c_{j}K$ is an integer for all $j$.
-Hence, by introducing new variables, we may replace the constraints $x_{j}=2^{c_{j}K}$ by the corresponding LMIs of small representation size from Theorem 5.3.
+Hence, by introducing new variables, we may replace the constraints $x_{j}=2^{c_{j}K}$ by the corresponding LMIs of small representation size from Theorem˜5.5.
 The resulting SDP can be computed in polynomial time.
 Furthermore, it is feasible if and only if the original SDP is feasible, which proves the statement.
-∎
 Corollary 5.9.
 There is a polynomial-time reduction from stopping simple stochastic games to the feasibility problem for semidefinite programs.
-Proof.
-Combine Corollary 5.4 with Proposition 3.2. ∎
-Acknowledgments
-This research benefited from the support of the FMJH Program PGMO.
-Manuel Bodirsky received funding from the ERC (Grant Agreement no.
-101071674, POCOCOP).
-Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency.
+Proof 5.10.
+We now present a strengthening of Corollary 5.7 which generalises it from max-average to max-plus-average CSPs.
+Theorem 5.11.
+There is a polynomial-time reduction from the max-plus-average constraint satisfaction problem to the feasibility problem for semidefinite programs.
+Proof 5.12.
+For any integer $k$, the singleton set $\{2^{k}\}$ can be efficiently represented as a projection of a spectrahedron (Theorem 5.5). Suppose that this spectrahedron is given by
+$Q^{(0)}+x_{1}Q^{(1)}+\dots+x_{n}Q^{(n)}\succeq 0,$
+and that the projection on the last coordinate $x_{n}$ is the singleton set $\{2^{k}\}$. Consider now the following spectrahedron, with one more variable $x_{0}$:
+$x_{0}Q^{(0)}+x_{1}Q^{(1)}+\dots+x_{n}Q^{(n)}\succeq 0\wedge x_{0}\geq 0.$
+We claim that the projection of this spectrahedron to the coordinates $(x_{0},x_{n})$ is precisely the closed half-line defined by
+$x_{n}=2^{k}x_{0}\wedge x_{0}\geq 0.$
+The proof has two steps. First, consider the projection to $(x_{0},x_{n})$ of the subset defined by
+$x_{0}Q^{(0)}+x_{1}Q^{(1)}+\dots+x_{n}Q^{(n)}\succeq 0\wedge x_{0}>0.$
+This projection is the open half-line defined by $x_{n}=2^{k}x_{0}\wedge x_{0}>0$.
+Indeed, take a point $(x_{0},x_{n})$ such that $x_{0}>0$.
+Moreover, suppose that we can find $(x_{1},\dots,x_{n-1})$ such that the matrix $x_{0}Q^{(0)}+x_{1}Q^{(1)}+\dots+x_{n}Q^{(n)}$ is positive semidefinite.
+Then the matrix
+$\frac{1}{x_{0}}(x_{0}Q^{(0)}+x_{1}Q^{(1)}+\dots+x_{n}Q^{(n)})=Q^{(0)}+\frac{x_{1}}{x_{0}}Q^{(1)}+\dots+\frac{x_{n}}{x_{0}}Q^{(n)}$
+is also positive semidefinite (here we use the assumption that $x_{0}>0$).
+Therefore, we get $\frac{x_{n}}{x_{0}}=2^{k}$ by the definition of the matrices $Q^{(i)}$.
+In particular, the projection is included in the half-line.
+Conversely, if $(x_{0},x_{n})$ are on the open half-line, then $\frac{x_{n}}{x_{0}}=2^{k}$ and the definition of the matrices $Q^{(i)}$ implies that there exist $(\tilde{x}_{1},\dots,\tilde{x}_{n})$ such that $Q^{(0)}+\tilde{x}_{1}Q^{(1)}+\dots+\tilde{x}_{n-1}Q^{(n-1)}+\frac{x_{n}}{x_{0}}Q^{(n)}\succeq 0$.
+Hence $x_{0}Q^{(0)}+x_{0}\tilde{x}_{1}Q^{(1)}+\dots+x_{0}\tilde{x}_{n-1}Q^{(n-1)}+x_{n}Q^{(n)}\succeq 0$ and $(x_{0},x_{n})$ belong to the projection.
+Second, since the projection of a spectrahedron is convex, the projection of
+$x_{0}Q^{(0)}+x_{1}Q^{(1)}+\dots+x_{n}Q^{(n)}\succeq 0\wedge x_{0}\geq 0$
+is either the open half-line or the closed half-line (we cannot add to the projection any other point $(x_{0},x_{n})$ with $x_{0}=0$ without destroying the convexity of the projection).
+Since $(0,0)$ is in the projection, we get that the projection is exactly the closed half-line as claimed.
+We have thus expressed the constraints of the form $x_{0}=2^{k}x_{1}$ as small SDPs.
+Since the constraint $x_{0}\leq 2^{k}x_{1}$ can be simulated by two constraints of the form $x_{0}\leq y$, $y=2^{k}x_{1}$, we can also express constraints of the form $x_{0}\leq 2^{k}x_{1}$ as small SDPs.
+By using the same reasoning as in the proof of Corollary 5.7, this allows us to simulate the non-Archimedean constraints of the form $x_{0}\leq t^{c}x_{1}$, and thus max-plus-average CSPs by Theorem˜4.1.